# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v4

@@ 1 Introduction
Influence maximization (IM) is a canonical network science problem where the goal is to select a small set of seed nodes in order to maximize the influence spread on a graph (Kempe et al.,, 2003; Chen et al.,, 2009). Over the past two decades, IM has garnered significant research interest due to its relevance in online marketing campaigns (Domingos and Richardson,, 2001; Bharathi et al.,, 2007; Chen et al., 2010b, ), rumor spreading (Kempe et al.,, 2003; Murata and Koga,, 2018), public health campaigns (Wilder et al.,, 2017; Yadav et al.,, 2017; Wilder et al.,, 2018; Yadav et al.,, 2018), vaccine strategies (Holme,, 2004; Lee et al.,, 2012), and more. Traditionally, the problem assumes the graph to be static, i.e., nodes and edges are fixed. In many situations, however, this may be unrealistic. For example, consider a marketing campaign on X where a company promotes a new product. As users and followers change daily, an effective IM algorithm must account for this dynamism. Therefore, recent work has looked at the IM problem on temporal networks (Holme and Saramäki,, 2012) where edges change with time. Since this combinatorial optimization problem is NP-hard, heuristics must be employed. Some methods use a greedy algorithm and probabilistic approximation of the expected influence spread (Aggarwal et al.,, 2012; Osawa and Murata,, 2015; Erkol et al.,, 2020), while others eschew costly influence spread calculations by simply ranking nodes in order of importance (Michalski et al.,, 2014; Murata and Koga,, 2018; Michalski et al.,, 2020). We refer the interested reader to Yanchenko et al., (2024) for a review of temporal IM.
@@
While IM is a well-known problem in the network, information and computer sciences, it has received relatively little attention from statisticians. This has led to several consequences on IM methodology development, not least of which being a lack of uncertainty quantification. Additionally, as the IM task is an optimization problem, statistics has a whole suit of techniques to offer that hitherto have been unexplored. One such approach is Bayesian Optimization (BO) which has emerged as a popular paradigm for optimizing complicated objective functions (Snoek et al.,, 2012; Shahriari et al.,, 2015; Frazier,, 2018). Machine learning hyper-parameter tuning (Snoek et al.,, 2012; Wu et al.,, 2019; Turner et al.,, 2021), experimental design (Frazier and Wang,, 2015; Shahriari et al.,, 2015; Greenhill et al.,, 2020) and computer experiments (Sürer et al.,, 2023; Sauer et al.,, 2023) are just a few of the many applications where BO has been successfully implemented. BO approximates a complicated objective function with a simple surrogate function and fits this model via Bayesian regression. The statistical model is maximized with an acquisition function before evaluating this new point on the original objective function. Finally, this point is added to the data set and the process repeats.
@@
In this work, we propose $\mathsf{BOPIM}$, Bayesian OPtimization for Influence Maximization, a BO algorithm for IM on temporal networks. The key challenges in applying a BO framework to our problem are in constructing the kernel function and optimizing the acquisition function. To address these, we propose two kernels, one based on the Hamming distance between seed sets, and the other on the similarity in node’s neighbors. We then leverage the Expected Improvement acquisition function to select the next candidate point, appropriately accounting for the noise in our observations. We optimize this function using a greedy algorithm which shares many similarities to the Trust Region framework (e.g., Eriksson et al.,, 2019; Wan et al.,, 2021; Papenmeier et al.,, 2023), but also accounts for the cardinality constraint. In numerical experiments on real-world networks, $\mathsf{BOPIM}$ yields seed sets with influence spreads comparable to those of a “gold-standard” greedy algorithm but at a fraction of the computational time.
@@
There are two primary contributions of this work. First, to the knowledge of the author, this is that first time that a BO framework has been applied to the IM problem. The key challenges stem from the inputs having a cardinality constraint, while also corresponding to non-Euclidean space, i.e., nodes on a graph. There have been several contributions to the combinatorial BO literature, including Garnett et al., (2010); Baptista and Poloczek, (2018); Oh et al., (2019); Ru et al., (2020); Buathong et al., (2020); Daulton et al., (2022); Papenmeier et al., (2023). The proposed method is inspired by these papers, most notably in the construction of the kernel, and optimization of the acquisition function. The cardinality constraint, however, means that these methods cannot easily be applied directly. For example, we show in detail why $\mathsf{COMBO}$ (Oh et al.,, 2019) cannot be used with such a constraint. Thus, this work addresses the various challenges of adapting the BO framework to the IM problem.
@@ 1.1 Notation and problem statement
We first must choose an influence diffusion mechanism, though the proposed method can easily be adapted to any diffusion mechanism. Some of the most popular choices are the Independent Cascade (IC) (Wang et al.,, 2012; Wen et al.,, 2017) and Linear Threshold (Chen et al., 2010a, ; Goyal et al.,, 2011) models. We adopt the Susceptible-Infected (SI) model from epidemiological studies (Osawa and Murata,, 2015; Murata and Koga,, 2018). In the SI model, a node is either in a susceptible ($\mathscr{S}$) or infected ($\mathscr{I}$) state. A node is infected if it was influenced at some point during the process, and susceptible otherwise. At time $t=1$, nodes in the seed set are set to state $\mathscr{I}$, and all other nodes initialize to state $\mathscr{S}$. At time $t$, a node in state $\mathscr{I}$ attempts to influence all of its neighbors. Specifically, if node $i$ is in state $\mathscr{I}$, node $j$ is in state $\mathscr{S}$, and there exists an edge between node $i$ and $j$ at time $t$, then node $j$ will be in state $\mathscr{I}$ at time $t+1$ with probability $\lambda$. The process concludes at time $T+1$ and the number of nodes in state $\mathscr{I}$ corresponds to the total influence spread. If we let $\sigma(S)$ be the expected number of influenced nodes at time $T+1$ for seed set $S$, the IM problem seeks the set of nodes of size $k$ that maximizes $\sigma(S)$ on $\mathcal{G}$, i.e.,
@@
Finding $S^{*}$ is a combinatorial optimization problem with a cardinality constraint and has been shown to be NP-hard under many popular diffusion mechanisms (Li et al.,, 2018) so finding the global optimum by brute-force calculation is infeasible even for moderate $n$. Indeed, as the influence spread function is typically computed using Monte Carlo (MC) simulations, even evaluating $\sigma(S)$ is costly. Many IM algorithms efficiently compute the influence spread function and apply a greedy algorithm to choose the seed nodes, while others skip evaluation of the objective function and instead use a heuristic to rank nodes by influence. In the remainder of this work, we describe a BO algorithm to approximate the solution in (1).
@@ 2.3.2 Kernel function
Recently, $\mathsf{COMBO}$ was proposed as an efficient and scalable method for combinatorial Bayesian optimization (Oh et al.,, 2019). This paper proposes a novel kernel function for combinatorial search spaces based on the idea of a combinatorial graph, where each vertex [^3] To attempt to avoid confusion between the combinatorial graph and the original graph where the information diffusion occurs, we will refer to the former as being made up of vertices, and the latter as possessing nodes. corresponds to a possible input to the objective function, and there is an edge between vertices if the inputs differ by only a single variable. The authors construct such a graph using the graph Cartesian product and derive a kernel based on the Graph Fourier Transform.
@@
In Oh et al., (2019), the authors show that the shortest distance between nodes on the combinatorial graph is equivalent to the Hamming distance. While we showed that $\mathsf{COMBO}$ cannot be used in our problem, we can borrow these ideas to construct a kernel function based on the Hamming distance. Specifically, let $\bm{x}_{i}$ and $\bm{x}_{j}$ be two seed sets. Then the covariance matrix is defined by
@@
where $\mathbb{I}(\cdot)$ is the indicator function. Because we are restricted to $k$ nodes in the seed set, if $k\leq n/2$, then $0\leq d_{H}\bm{x}_{i},\bm{x}_{j})\leq 2k$, which is why we divide by $2k$ instead of $n$. This kernel is in a similar spirit to those of $\mathsf{CASMOPOLITAN}$ (Wan et al.,, 2021), $\mathsf{Bounce}$ (Papenmeier et al.,, 2023) and $\mathsf{CoCaBO}$ (Ru et al.,, 2020). Please see the Supplemental Materials for a proof that this kernel is positive semi-definite.
@@
Recall that for the IM task, we initially influence a small set of nodes which then propagate information to the rest of the network. Clearly, these initially selected nodes (as well as their neighboring nodes) are very important for the final influence spread (Erkol et al.,, 2020). Indeed, if the nodes corresponding to seed sets $\bm{x}_{i}$ and $\bm{x}_{j}$ have many neighbors in common, then we would expect their influence spreads also to be similar.
Inspired by this observation, as well as Garnett et al., (2010), who studies the BO task on combinatorial search-spaces, we propose the following kernel function. Let $\{v_{i_{1}},\dots,v_{i_{k}}\}$ correspond to seed set $\bm{x}_{i}$. First, we find all neighbors of $\{v_{i_{1}},\dots,v_{i_{k}}\}$ at time $t=1$, and call this neighboring set $\mathcal{S}_{i}$. In other words, if node $v_{\ell}$ has an incoming edge from some node in $\{v_{i_{1}},\dots,v_{i_{k}}\}$ at time $t=1$, then $v_{\ell}\in\mathcal{S}_{i}$. Of course, $\{v_{i_{1}},\dots,v_{i_{k}}\}\in\mathcal{S}_{i}$ as well. We also compute the set of neighboring nodes set, $\mathcal{S}_{j}$, for $\bm{x}_{j}$. Again, if $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ have many common elements, then we want $\kappa(\bm{x}_{i},\bm{x}_{j})$ to be large. Therefore, we can compute the Jaccard coefficient between $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ and use this as our kernel function. The Jaccard coefficient (JC) (Jaccard,, 1901, 1912) is a simple metric to quantify the similarity between two finite sets. If $A$ and $B$ are two such sets, then the JC is the number of common elements between the two sets divided by the total number of elements in each set, i.e.,
@@
The use of the JC in such context enjoys some precedence. First, the JC has been used for link prediction in graphs, most notably in (Liben-Nowell and Kleinberg,, 2003). Additionally, Gosnell and Evangelou, (2024) use the Tanimoto index, another name for the JC, to construct a GP kernel with drug discovery applications. More generally, the JC has received significant attention in chemoinformatics (Rogers and Tanimoto,, 1960; Gower,, 1971; Todeschini et al.,, 2012; Moss and Griffiths,, 2020; Gosnell and Evangelou,, 2024). Finally, the JC (Tanimoto index) has been shown to yield a positive semi-definite correlation matrix, making it a valid choice for our kernel function (Gower,, 1971; Bouchard et al.,, 2013).
@@ 2.4 Acquisition function
We prefer the Expected Improvement acquisition function (Močkus,, 1975; Jones et al.,, 1998; Frazier,, 2018) as it is one of the most popular acquisition functions and does not require any hyper-parameter tuning. The idea is to evaluate the objective function at the value which leads to the largest expected increase compared to the current maximum. Specifically, assume that we are in $b$-th iteration of the BO loop. Let $\bm{x}\in\{0,1\}^{n}$ and $\tilde{\mu}^{(b)}(\bm{x})$ and ${\tilde{\sigma}}^{(b)}(\bm{x})$ be the conditional mean and standard deviation, respectively, of our GP model given the current data, where we explicitly emphasize the dependence on the iteration with the super-script in the notation. By properties of the multivariate normal distribution, we have
$\tilde{\mu}^{(b)}(\bm{x})=\beta_{0}^{(b)}{\bf 1}_{N_{0}+b}+\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-\beta_{0}^{(b)}{\bf 1}_{N_{0}+b})$
@@
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function and distribution function, respectively, of the standard normal (Frazier,, 2018). Recall, however, that the influence spread, $y_{i}$, is observed with Gaussian noise $\varepsilon_{i}$, making ${f^{(b)}}^{*}$ a (potentially) poor and non-robust estimate of the global maximum, particularly if it was observed with large noise. Following Vazquez et al., (2008) and Picheny et al., (2013) , we replace ${f^{(b)}}^{*}$ with a “plug-in” estimate of the global maximum; namely, the maximum of the conditional mean function, $\tilde{\mu}^{(b)}(\cdot)$, evaluated at the previously sampled points, i.e.,
@@
We are still not quite finished as the acquisition function in (5) implies that we observe the next observation deterministically. To account for the noise in the ensuing observations, we included a multiplicative term as in Huang et al., 2006a ; Huang et al., 2006b , which penalizes sampling a new seed set where the conditional standard deviation, ${\tilde{\sigma}}^{(b)}(\cdot)$ is small. This is a sensible choice, as a small variance implies we already know the influence spread of this seed set with fairly high certainty, encouraging exploring elsewhere in the input space. The final (augmented) expected improvement (AEI) acquisition function becomes
@@
The astute reader will notice that we have simply traded one combinatorial optimization problem (maximizing $f(\cdot)$ for maximizing $AEI(\cdot)$). $AEI(\cdot)$, however, is much cheaper to evaluate than $f(\cdot)$ so finding its optimum will be significantly faster. Indeed, we propose a simple greedy algorithm to find the maximum of $AEI(\cdot)$ with details in the Supplemental Materials. The main idea of this algorithm is that nodes in the current seed set are swapped with other nodes. If the switch improves the solution (larger $AEI(\cdot)$), then its kept, otherwise the previous seed set is used. In this way, the algorithm makes the locally optimal decision. The algorithm continues until all nodes have been looped through and no changes are made. Please see the Supplemental Materials for a connection between the proposed greedy algorithm and well-known Trust region (TR) framework (e.g., Eriksson et al.,, 2019; Wan et al.,, 2021; Papenmeier et al.,, 2023)
@@
       Find $\tilde{\bm{x}}=\arg\max_{\bm{x},\sum_{j}x_{j}=k}AEI(\bm{x})$   Compute $y=f(\tilde{\bm{x}})$  Append observation and re-compute posterior of $\beta_{0}$ and $\sigma^{2}$  Update ${\bf K}$

### v5

@@ 1 Introduction
Influence maximization (IM) is a canonical network science problem where the goal is to select a small set of seed nodes in order to maximize the influence spread on a graph (Kempe et al., 2003; Chen et al., 2009). Over the past two decades, IM has garnered significant research interest due to its relevance in online marketing campaigns (Domingos and Richardson, 2001; Bharathi et al., 2007; Chen et al., 2010b), rumor spreading (Kempe et al., 2003; Murata and Koga, 2018), public health campaigns (Wilder et al., 2017; Yadav et al., 2017; Wilder et al., 2018; Yadav et al., 2018), vaccine strategies (Holme, 2004; Lee et al., 2012), and more. Traditionally, the problem assumes the graph to be static, i.e., nodes and edges are fixed. In many situations, however, this may be unrealistic. For example, consider a marketing campaign on X where a company promotes a new product. As users and followers change daily, an effective IM algorithm must account for this dynamism. Therefore, recent work has looked at the IM problem on temporal networks (Holme and Saramäki, 2012) where edges change with time. Since this combinatorial optimization problem is NP-hard, heuristics must be employed. Some methods use a greedy algorithm and probabilistic approximation of the expected influence spread (Aggarwal et al., 2012; Osawa and Murata, 2015; Erkol et al., 2020), while others eschew costly influence spread calculations by simply ranking nodes in order of importance (Michalski et al., 2014; Murata and Koga, 2018; Michalski et al., 2020). We refer the interested reader to Yanchenko et al. (2024) for a review of temporal IM.
@@
While IM is a well-known problem in the network, information and computer sciences, it has received relatively little attention from statisticians. This has led to several consequences on IM methodology development, not least of which being a lack of uncertainty quantification. Additionally, as the IM task is an optimization problem, statistics has a whole suit of techniques to offer that hitherto have been unexplored. One such approach is Bayesian Optimization (BO) which has emerged as a popular paradigm for optimizing complicated objective functions (Snoek et al., 2012; Shahriari et al., 2015; Frazier, 2018). Machine learning hyper-parameter tuning (Snoek et al., 2012; Wu et al., 2019; Turner et al., 2021), experimental design (Frazier and Wang, 2015; Shahriari et al., 2015; Greenhill et al., 2020) and computer experiments (Sürer et al., 2023; Sauer et al., 2023) are just a few of the many applications where BO has been successfully implemented. BO approximates a complicated objective function with a simple surrogate function and fits this model via Bayesian regression. The statistical model is maximized with an acquisition function before evaluating this new point on the original objective function. Finally, this point is added to the data set and the process repeats.
@@
In this work, we propose $\mathsf{BOPIM}$, Bayesian OPtimization for Influence Maximization, a BO algorithm for IM on temporal networks. The key challenges in applying a BO framework to our problem are in constructing the kernel function and optimizing the acquisition function. To address these, we propose two kernels, one based on the Hamming distance between seed sets, and the other on the similarity in node’s neighbors. We then leverage the Expected Improvement acquisition function to select the next candidate point, appropriately accounting for the noise in our observations. We optimize this function using a greedy algorithm which shares many similarities to the Trust Region framework (e.g., Eriksson et al., 2019; Wan et al., 2021; Papenmeier et al., 2023), but also accounts for the cardinality constraint. In numerical experiments on real-world networks, $\mathsf{BOPIM}$ yields seed sets with influence spreads comparable to those of a “gold-standard” greedy algorithm but at a fraction of the computational time.
@@
There are two primary contributions of this work. First, to the knowledge of the author, this is that first time that a BO framework has been applied to the IM problem. The key challenges stem from the inputs having a cardinality constraint, while also corresponding to non-Euclidean space, i.e., nodes on a graph. There have been several contributions to the combinatorial BO literature, including Garnett et al. (2010); Baptista and Poloczek (2018); Oh et al. (2019); Ru et al. (2020); Buathong et al. (2020); Daulton et al. (2022); Papenmeier et al. (2023). The proposed method is inspired by these papers, most notably in the construction of the kernel, and optimization of the acquisition function. The cardinality constraint, however, means that these methods cannot easily be applied directly. For example, we show in detail why $\mathsf{COMBO}$ (Oh et al., 2019) cannot be used with such a constraint. Thus, this work addresses the various challenges of adapting the BO framework to the IM problem.
@@ 1.1 Notation and problem statement
We first must choose an influence diffusion mechanism, though the proposed method can easily be adapted to any diffusion mechanism. Some of the most popular choices are the Independent Cascade (IC) (Wang et al., 2012; Wen et al., 2017) and Linear Threshold (Chen et al., 2010a; Goyal et al., 2011) models. We adopt the Susceptible-Infected (SI) model from epidemiological studies (Osawa and Murata, 2015; Murata and Koga, 2018). In the SI model, a node is either in a susceptible ($\mathscr{S}$) or infected ($\mathscr{I}$) state. A node is infected if it was influenced at some point during the process, and susceptible otherwise. At time $t=1$, nodes in the seed set are set to state $\mathscr{I}$, and all other nodes initialize to state $\mathscr{S}$. At time $t$, a node in state $\mathscr{I}$ attempts to influence all of its neighbors. Specifically, if node $i$ is in state $\mathscr{I}$, node $j$ is in state $\mathscr{S}$, and there exists an edge between node $i$ and $j$ at time $t$, then node $j$ will be in state $\mathscr{I}$ at time $t+1$ with probability $\lambda$. The process concludes at time $T+1$ and the number of nodes in state $\mathscr{I}$ corresponds to the total influence spread. If we let $\sigma(S)$ be the expected number of influenced nodes at time $T+1$ for seed set $S$, the IM problem seeks the set of nodes of size $k$ that maximizes $\sigma(S)$ on $\mathcal{G}$, i.e.,
@@
Finding $S^{*}$ is a combinatorial optimization problem with a cardinality constraint and has been shown to be NP-hard under many popular diffusion mechanisms (Li et al., 2018) so finding the global optimum by brute-force calculation is infeasible even for moderate $n$. Indeed, as the influence spread function is typically computed using Monte Carlo (MC) simulations, even evaluating $\sigma(S)$ is costly. Many IM algorithms efficiently compute the influence spread function and apply a greedy algorithm to choose the seed nodes, while others skip evaluation of the objective function and instead use a heuristic to rank nodes by influence. In the remainder of this work, we describe a BO algorithm to approximate the solution in (1).
@@ 2.3.2 Kernel function
Recently, $\mathsf{COMBO}$ was proposed as an efficient and scalable method for combinatorial Bayesian optimization (Oh et al., 2019). This paper proposes a novel kernel function for combinatorial search spaces based on the idea of a combinatorial graph, where each vertex [^3] To attempt to avoid confusion between the combinatorial graph and the original graph where the information diffusion occurs, we will refer to the former as being made up of vertices, and the latter as possessing nodes. corresponds to a possible input to the objective function, and there is an edge between vertices if the inputs differ by only a single variable. The authors construct such a graph using the graph Cartesian product and derive a kernel based on the Graph Fourier Transform.
@@
In Oh et al. (2019), the authors show that the shortest distance between nodes on the combinatorial graph is equivalent to the Hamming distance. While we showed that $\mathsf{COMBO}$ cannot be used in our problem, we can borrow these ideas to construct a kernel function based on the Hamming distance. Specifically, let $\bm{x}_{i}$ and $\bm{x}_{j}$ be two seed sets. Then the covariance matrix is defined by
@@
where $\mathbb{I}(\cdot)$ is the indicator function. Because we are restricted to $k$ nodes in the seed set, if $k\leq n/2$, then $0\leq d_{H}\bm{x}_{i},\bm{x}_{j})\leq 2k$, which is why we divide by $2k$ instead of $n$. This kernel is in a similar spirit to those of $\mathsf{CASMOPOLITAN}$ (Wan et al., 2021), $\mathsf{Bounce}$ (Papenmeier et al., 2023) and $\mathsf{CoCaBO}$ (Ru et al., 2020). Please see the Supplemental Materials for a proof that this kernel is positive semi-definite.
@@
Recall that for the IM task, we initially influence a small set of nodes which then propagate information to the rest of the network. Clearly, these initially selected nodes (as well as their neighboring nodes) are very important for the final influence spread (Erkol et al., 2020). Indeed, if the nodes corresponding to seed sets $\bm{x}_{i}$ and $\bm{x}_{j}$ have many neighbors in common, then we would expect their influence spreads also to be similar.
Inspired by this observation, as well as Garnett et al. (2010), who studies the BO task on combinatorial search-spaces, we propose the following kernel function. Let $\{v_{i_{1}},\dots,v_{i_{k}}\}$ correspond to seed set $\bm{x}_{i}$. First, we find all neighbors of $\{v_{i_{1}},\dots,v_{i_{k}}\}$ at time $t=1$, and call this neighboring set $\mathcal{S}_{i}$. In other words, if node $v_{\ell}$ has an incoming edge from some node in $\{v_{i_{1}},\dots,v_{i_{k}}\}$ at time $t=1$, then $v_{\ell}\in\mathcal{S}_{i}$. Of course, $\{v_{i_{1}},\dots,v_{i_{k}}\}\in\mathcal{S}_{i}$ as well. We also compute the set of neighboring nodes set, $\mathcal{S}_{j}$, for $\bm{x}_{j}$. Again, if $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ have many common elements, then we want $\kappa(\bm{x}_{i},\bm{x}_{j})$ to be large. Therefore, we can compute the Jaccard coefficient between $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$ and use this as our kernel function. The Jaccard coefficient (JC) (Jaccard, 1901, 1912) is a simple metric to quantify the similarity between two finite sets. If $A$ and $B$ are two such sets, then the JC is the number of common elements between the two sets divided by the total number of elements in each set, i.e.,
@@
The use of the JC in such context enjoys some precedence. First, the JC has been used for link prediction in graphs, most notably in (Liben-Nowell and Kleinberg, 2003). Additionally, Gosnell and Evangelou (2024) use the Tanimoto index, another name for the JC, to construct a GP kernel with drug discovery applications. More generally, the JC has received significant attention in chemoinformatics (Rogers and Tanimoto, 1960; Gower, 1971; Todeschini et al., 2012; Moss and Griffiths, 2020; Gosnell and Evangelou, 2024). Finally, the JC (Tanimoto index) has been shown to yield a positive semi-definite correlation matrix, making it a valid choice for our kernel function (Gower, 1971; Bouchard et al., 2013).
@@ 2.4 Acquisition function
We prefer the Expected Improvement acquisition function (Močkus, 1975; Jones et al., 1998; Frazier, 2018) as it is one of the most popular acquisition functions and does not require any hyper-parameter tuning. The idea is to evaluate the objective function at the value which leads to the largest expected increase compared to the current maximum. Specifically, assume that we are in $b$-th iteration of the BO loop. Let $\bm{x}\in\{0,1\}^{n}$ and $\tilde{\mu}^{(b)}(\bm{x})$ and ${\tilde{\sigma}}^{(b)}(\bm{x})$ be the conditional mean and standard deviation, respectively, of our GP model given the current data, where we explicitly emphasize the dependence on the iteration with the super-script in the notation. By properties of the multivariate normal distribution, we have
$\tilde{\mu}^{(b)}(\bm{x})=\beta_{0}^{(b)}+\kappa^{*}(\bm{x},{\bf X})^{T}(\gamma{\bf K}+{\bf I}_{N_{0}+b})^{-1}(\bm{y}-\beta_{0}^{(b)}{\bf 1}_{N_{0}+b})$
@@
where $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function and distribution function, respectively, of the standard normal (Frazier, 2018). Recall, however, that the influence spread, $y_{i}$, is observed with Gaussian noise $\varepsilon_{i}$, making $f^{(b)}$$^{*}$ a (potentially) poor and non-robust estimate of the global maximum, particularly if it was observed with large noise. Following Vazquez et al. (2008) and Picheny et al. (2013) , we replace $f^{(b)}$$^{*}$ with a “plug-in” estimate of the global maximum; namely, the maximum of the conditional mean function, $\tilde{\mu}^{(b)}(\cdot)$, evaluated at the previously sampled points, i.e.,
@@
We are still not quite finished as the acquisition function in (5) implies that we observe the next observation deterministically. To account for the noise in the ensuing observations, we included a multiplicative term as in Huang et al. (2006a, b), which penalizes sampling a new seed set where the conditional standard deviation, ${\tilde{\sigma}}^{(b)}(\cdot)$ is small. This is a sensible choice, as a small variance implies we already know the influence spread of this seed set with fairly high certainty, encouraging exploring elsewhere in the input space. The final (augmented) expected improvement (AEI) acquisition function becomes
@@
The astute reader will notice that we have simply traded one combinatorial optimization problem (maximizing $f(\cdot)$ for maximizing $AEI(\cdot)$). $AEI(\cdot)$, however, is much cheaper to evaluate than $f(\cdot)$ so finding its optimum will be significantly faster. Indeed, we propose a simple greedy algorithm to find the maximum of $AEI(\cdot)$ with details in the Supplemental Materials. The main idea of this algorithm is that nodes in the current seed set are swapped with other nodes. If the switch improves the solution (larger $AEI(\cdot)$), then its kept, otherwise the previous seed set is used. In this way, the algorithm makes the locally optimal decision. The algorithm continues until all nodes have been looped through and no changes are made. Please see the Supplemental Materials for a connection between the proposed greedy algorithm and well-known Trust region (TR) framework (e.g., Eriksson et al., 2019; Wan et al., 2021; Papenmeier et al., 2023)
@@
[…]